Circular No. 6462

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

1996 PW M. Hammergren, Department of Astronomy, University of Washing-

ton, reports: “Low-resolution reflectance spectra (range 400–1000 nm) of the unusual object 1996 PW were obtained on Aug. 19.42 UT with the Apache Point Observatory 3.5-m telescope (+ DIS). The spectra display a featureless red-sloped continuum with a reflectivity gradient S′ (Luu 1993, Icarus 104, 138) of 10.2 ± 0.4 percent (1000 Å)^{−1} over the wavelength range 620–860 nm. The object appears significantly redder than the colors reported on IAUC 6452, which are bluer than solar. No cometary emission features are seen; however, these spectral characteristics are consistent with those of cometary nuclei and D-type asteroids (Luu, op.cit.; Jewitt and Luu 1990, A.J. 100, 933).”

CYGNUS X-1 S. N. Zhang, B. A. Harmon, W. S. Paciesas, C. R. Robinson, D. M.

Scott, and W. Yu, on behalf of the BATSE/CGRO team; and R. Remillard and W. Cui, on behalf of the ASM/RXTE team, report: “The blackhole candidate Cyg X-1 is returning to the hard state, following the recent hard x-ray flux increase (IAUC 6447). Hard x-ray (20–200 keV) and soft x-ray (1.3–12 keV) fluxes are very close to the values before the hard-to-soft-state transition in mid-May 1995 (IAUC 6405). Spectral hardening has also been observed in both bands. The ASM (1.3–12 keV) and BATSE (20–200 keV) long-term lightcurves are anticorrelated, with the total 1.3- to 200-keV luminosity remaining approximately unchanged within 15 percent during the entire state transition from mid-May to late August 1996. The BATSE and ASM soft-state power-density spectra (PDS) of Cyg X-1 (roughly 10^{−7}– 10^{−4} Hz), which can be approximately represented by a power law with an index of ∼ −1.0, are apparently a low-frequency extension of the same softstate PDS detected by PCA/RXTE above 10^{−3} Hz (Cui et al., submitted to Ap.J. Lett.). Flaring episodes separated by ∼ 45–50 days and of duration 5–30 days were observed superimposed on the spectral transition in both ASM and BATSE data. (The mid-1991 to early-1996 BATSE lightcurve for Cyg X-1 does not show evidence of any significant periodicity near 50 days.) A possible 15-day delay of the flares in the ASM band with respect to the flares in the BATSE band was also detected from the cross correlation between ASM and BATSE lightcurves.”

1996 August 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green